# On the Optimality of Coded Distributed Computing for Ring Networks

### v1

@@ III Main Results
(Uncoded Schemes) Follow the same steps in the coded schemes in Sections IV and V, except that each node directly forwarding each intermediate value instead of generating coded messages, then the NCL of the uncoded schemes for all-gather and all-to-all are $\mathsf{T}^{\textnormal{uncoded}}_{1}(r,d)=2\mathsf{T}_{1}(r,d)$ and $\mathsf{T}^{\textnormal{uncoded}}_{\textnormal{2-cyc}}(r,d)=2\mathsf{T}^{\textnormal{ach}}_{\textnormal{2-cyc}}(r,d),$ respectively. This indicates that ring topology would constrain the multicast opportunities, i.e., the proposed coded scheme can at most offer a multiplicative factor of $2$ of benefit compared with the uncoded scheme. Notably, prior work in network coding has shown that coding gain is bounded by a factor of $2$ for the information transmission in various undirected network model [32, 33], while they do not consider the redundant computation (which can be modeled by allowing the same source messages to be sent from multiple nodes) and assume the network only has a single communication session.
Figure 5: Comparison of upper bound and lower bound of all-gather.
Figure 6: Comparison of upper bound and lower bound of all-to-all.
@@
(Advantanges of Parallelism) An advantage of our proposed schemes is their potential for parallelism. For example, when $d=1$, the nodes in $\{n_{i},n_{i+2},n_{i+4},\ldots\}$ can transmit messages concurrently during a given time step without interfering with each other, and the nodes in $\{n_{i+1},n_{i+3},n_{i+5},\ldots\}$ receive their intended messages. In the next step, the roles of the sending and receiving nodes are reversed to send other messages. This enables a more efficient communication bandwidth utilization, compared to the work in [9] where nodes contend for access to a shared communication link.
@@ IV-A An Illustrative Example
Obviously, node $n_{i}$ can decode $V_{i-1}$ and $V_{i+2}$ from $V_{i-1}\oplus V_{i}$ and $V_{i+1}\oplus V_{i+2}$, as it already caches $V_{i}$ and $V_{i+1}$. Then, using the decoded symbols $V_{i-1}$ and $V_{i+2}$, node $n_{i}$ can then proceed to decode new symbols $V_{i-2}$ and $V_{i+3}$. This successive decoding process continues further for $V_{i-3}$ and $V_{i+4}$. In other words, each node decodes symbols sequentially, starting from those sent by nearby nodes and progressing to those from farther nodes. The decoding order is presented in Table I. The symbol ‘*’ at row $n_{i}$ and column $V_{j}$, where $i\in[8]$ and $j\in[8]$, indicates that node $n_{i}$ generates (and thus has) the IV $V_{j}$. In row $n_{i}$, integer $1$ indicates that the IV in the corresponding column can be directly decoded by node $n_{i}$ upon receiving the encoded symbol broadcast by its neighboring nodes. The integer $2$ signifies that decoding the corresponding IV depends on the symbol associated with the nearest integer $1$ in the same row. Similarly, the integer $3$ denotes that decoding the corresponding IV relies on the symbol associated with the nearest integer $2$ in the same row.
TABLE I: Decoding order for $\mathsf{N}=8$, $r=2$ and $d=3$
@@ IV-B General Scheme
If $d\geq r$, node $n_{i}$ can also decode $2d$ desired symbols but with successive decoding. Taking $Y_{i_{-}}^{(1)}$ as an example (illustrated in Fig. 8), the same color on the two sides of the summations means the different coded symbols share components. Since node $n_{i}$ has already known the right-hand side of the first $r-1$ summations (i.e., $V_{i+r-2},\ldots,V_{i}$), it can decode their left-hand side. These decoded symbols then serve as known components for the next $r-1$ summations (i.e., the $r^{\textnormal{th}}$ to $(2r-1)^{\textnormal{th}}$ summations from the bottom), allowing further decoding in a successive manner. Following this process, node $n_{i}$ successively decodes $2d$ desired symbols
@@
We now describe the transmission process from time clock $k\geq 2$. After time $k-1$, node $n_{i}$ has successfully decoded and stored the following symbols,
@@ V Coded Transmission for All-to-all Computing
For clarity, we re-define the notation for transmitted messages. The message $X_{i}^{(t)}$, transmitted by node $n_{i}$ at time slot $t$, is now written as $X_{i}^{(j,k)}$. Here, $X_{i}^{(j,k)}$ represents the message transmitted by node $n_{i}$ at the $k^{\textnormal{th}}$ step of round $j$, where $0\leq k\leq j$ and $t=1+2+\ldots+(j-1)+k$. Similarly, we re-write $Y_{i}^{(t)}$ as $Y_{i}^{(j,k)}$, maintaining the same relationship $0\leq k\leq j$ and $1+2+\ldots+(j-1)+k=t$.
@@ V-A An Illustrative Example
In the first round, each node transmits packets with distance $1$ to their destination nodes, i.e., $W_{i}^{(1)}$, $i\in[\mathsf{N}]$. The encoding and transmission process is presented in Table III. For each node $i$, the coded packet $X_{i}^{(1,1)}=v_{i}^{i+1}\oplus v_{i+2}^{i-1}$ is broadcast to neighbor nodes $n_{i-1}$ and $n_{i+1}$. For example, node $n_{1}$ broadcasts $X_{1}^{(1,1)}=v_{1}^{2}\oplus v_{3}^{8}$, it also receives the encoded packets $X_{2}^{(1,1)}=v_{2}^{3}\oplus v_{4}^{1}$ and $X_{8}^{(1,1)}=v_{8}^{1}\oplus v_{2}^{7}$ which are broadcast by node $n_{2}$ and $n_{8}$, respectively. Since node $n_{1}$ has cached $v_{2}^{7}$ and $v_{2}^{3}$, it can decode the desired packets $v_{8}^{1}$ and $v_{4}^{1}$ by subtracting the known values from $X_{2}^{(1,1)}=v_{2}^{3}\oplus v_{4}^{1}$ and $X_{8}^{(1,1)}=v_{8}^{1}\oplus v_{2}^{7}$, respectively. The same decoding process applies to all other nodes.
In the second round, it will take $2$ steps of each node to transmit the packets that have a distance of $2$, i.e., $W_{i}^{(2)}$, $i\in[\mathsf{N}]$. The encoding and transmission process is shown in Table III. For each node, the encoded packets of the first broadcast are generated from the packets in $W_{i}^{(2)}$. After receiving the coded packets from the broadcast in the first step, each node eliminates the cached content and re-encodes the remaining packets for the broadcast in the second step. For example, node $n_{1}$ received $X_{8}^{(2,1)}=v_{8}^{2}\oplus v_{2}^{6}$ from node $n_{8}$ and $X_{2}^{(2,1)}=v_{2}^{4}\oplus v_{4}^{8}$ from node $n_{2}$ in the first step. Node $n_{1}$ can decode $v_{8}^{2}$ and $v_{4}^{8}$ since it has cached $v_{2}^{6}$ and $v_{2}^{4}$. Then $n_{1}$ generates a new coded packet $X_{1}^{(2,2)}=v_{8}^{2}\oplus v_{4}^{8}$ and broadcasts it in the second step.
The third round requires $3$ steps to send the packets $W_{i}^{(3)}$, $i\in[\mathsf{N}]$, as shown in Tables IV. The coding process follows the same principle as in previous rounds. However, we can observe that in this case, each packet in $W_{i}^{(3)}$ will be sent by two nodes. For example, $v_{1}^{4}$ appears in both coded packets $X_{1}^{(3,1)}=v_{1}^{4}\oplus v_{3}^{6}$ and $X_{7}^{(3,1)}=v_{7}^{2}\oplus v_{1}^{4}$, broadcast by $n_{1}$ and $n_{7}$, respectively. Therefore, the two nodes can each send half of the packets to save on transmission costs. Finally, the NCL can be calculated as $\mathsf{T}_{2}=1+2+2.5=5.5$. In this example, the decoding at each step relies solely on initially cached content. However, as we will see in the next section, from the third step onward, decoding also depends on the undesired symbols decoded in the previous two steps.
TABLE II: Round $1$: Step $1$.
Round $1$ Step $1$
@@
TABLE III: Round $2$: Step $1$ and Step $2$.
Round $2$ Step $1$ Step $2$
@@
TABLE IV: Round $3$: Step $1$, step $2$ and Step $3$.
Round $3$ Step $1$ Step $2$ Step $3$
@@ V-B1 Proposed scheme for $d=1$
As described in Section V-A, the transmitted IVs of node $i$ can be categorized into different sets $W_{i}^{(1)},\ldots,W_{i}^{(\lceil\frac{\mathsf{N}-r}{2}\rceil)}$. Since the packets follow a cyclic placement, the desired packets are located at most $\left\lceil\frac{\mathsf{N}-r}{2}\right\rceil$ hops away from node $n_{i}$. The transmissions are performed in rounds, with packets at different distances being sent in separate rounds.
@@
$=O\left(\frac{\mathsf{N}}{4}(\frac{\mathsf{N}}{2}\!\!-\!\!r)\!+\!\frac{(\mathsf{N}-r)}{4}\!+\!\frac{r^{2}}{8}\right).$ (18)
@@ V-B2 Modification for $1<d\leq 2(r-1)$
The number of broadcasts can be further reduced when the nodes have a larger broadcast distance. Based on the transmission strategy for $d=1$, here is the modified version for $d\leq 2(r-1)$. Unlike the case of $d=1$ where the $3$ adjacent nodes perform reverse carpooling, here we allow nodes that are distance $d$ or $r-1$ apart to perform reverse carpooling. The paths of information flow are demonstrated in Fig. 10.
@@
$=\underbrace{1+\ldots+1}_{d_{1}}+\underbrace{2+\ldots+2}_{d}+\ldots+\!\underbrace{\left\lceil\frac{\frac{\mathsf{N}-r}{2}\!-\!d_{1}}{d}\right\rceil\!+\!1\!+\!\ldots\!+\!\left\lceil\frac{\frac{\mathsf{N}-r}{2}\!-\!d_{1}}{d}\right\rceil\!+\!1}_{\left(\lceil\frac{\mathsf{N}-r}{2}\rceil-d_{1}\right)_{\bmod d}}$
$=O\left(\frac{\mathsf{N}}{4d}(\frac{\mathsf{N}}{2}\!-\!r)\!+\!\frac{5(\mathsf{N}\!-\!r)}{4}\!+\!\frac{r^{2}\!-\!4d_{1}(\mathsf{N}\!-\!r\!+\!d\!-\!d_{1})}{2d}\right)$
=$=\!\left\{\begin{array}[]{cc}\vspace{5pt}\!\!\!\!O\left(\frac{\mathsf{N}}{4d}(\frac{\mathsf{N}}{2}\!\!-\!\!3r)\!+\!\frac{3\mathsf{N}-5r}{4}\!+\!\frac{9r^{2}+4(\mathsf{N}-r+1)}{8d}\right)&\begin{subarray}{c}d_{1}=r-1,\\ 1<r,\end{subarray}\\ \!\!\!\!O\left(\frac{\mathsf{N}}{4d}(\frac{\mathsf{N}}{2}\!\!-\!\!r)\!+\!\frac{(\mathsf{N}-r)}{4}\!+\!\frac{r^{2}}{8d}\right)&\begin{subarray}{c}d_{1}=d,\\ 1<r.\end{subarray}\end{array}\right.$ (24)
@@ V-B3 Modification for $2r-1\leq d$
to nodes $n_{i-d}$ and $n_{i+d}$, respectively. For $i\in[\mathsf{N}],j\in[\lceil\frac{\mathsf{N}-r}{2}\rceil]$, $X_{i}^{(j,0)}$ are additional transmission before the first step. The remaining steps of the process can proceed as described previously. Therefore, in this case, we can calculate the NCL as
@@
$\mathsf{T}_{\textnormal{2-cyc}}(r,d)$ $=\underbrace{1+\ldots+1}_{d}+\underbrace{2+\ldots+2}_{d}+\ldots+\!\underbrace{\left\lceil\frac{\frac{\mathsf{N}-r}{2}}{d}\right\rceil\!+\!\ldots\!+\!\left\lceil\frac{\frac{\mathsf{N}-r}{2}}{d}\right\rceil}_{\lceil\frac{\mathsf{N}-r}{2}\rceil_{\bmod d}}+\left\lceil\frac{\mathsf{N}-r}{2}\right\rceil$
@@
$=O\left(\frac{\mathsf{N}}{4d}(\frac{\mathsf{N}}{2}\!\!-\!\!r)\!+\!\frac{3(\mathsf{N}-r)}{4}\!+\!\frac{r^{2}}{8d}\right).$ (25)
@@
From (18), (24) and (25), we can conclude that the achievable NCL of the proposed scheme is as stated in Theorem 2.
VI Conclusion

### v2

@@ III Main Results
(Uncoded Schemes) Follow the same steps in the coded schemes in Sections IV and V, but letting each node directly forward each intermediate value instead of generating coded messages, the NCL of the uncoded schemes for all-gather and all-to-all are $\mathsf{T}^{\textnormal{uncoded}}_{1}(r,d)=2\mathsf{T}_{1}(r,d)$ and $\mathsf{T}^{\textnormal{uncoded}}_{\textnormal{2-cyc}}(r,d)=2\mathsf{T}^{\textnormal{ach}}_{\textnormal{2-cyc}}(r,d),$ respectively. This factor-of-two gap arises because local broadcasts without coding cannot simultaneously satisfy distinct demands flowing in opposite directions. This indicates that ring topology would constrain the multicast opportunities, i.e., the proposed coded scheme can at most offer a multiplicative factor of $2$ of benefit compared with the uncoded scheme. Notably, prior work in network coding has shown that coding gain is bounded by a factor of $2$ for the information transmission in various undirected network model [36, 37], while they do not consider the redundant computation (which can be modeled by allowing the same source messages to be sent from multiple nodes) and assume the network only has a single communication session.
@@
(Advantages of Parallelism) An advantage of our proposed schemes is their potential for parallelism. For example, when $d=1$, the nodes in $\{n_{i},n_{i+2},n_{i+4},\ldots\}$ can transmit messages concurrently during a given time step without interfering with each other, and the nodes in $\{n_{i+1},n_{i+3},n_{i+5},\ldots\}$ receive their intended messages. In the next step, the roles of the sending and receiving nodes are reversed to send other messages. This enables a more efficient communication bandwidth utilization, compared to the work in [9] where nodes contend for access to a shared communication link.
@@ IV-A An Illustrative Example
Obviously, node $n_{i}$ can decode $V_{i-1}$ and $V_{i+2}$ from $V_{i-1}\oplus V_{i}$ and $V_{i+1}\oplus V_{i+2}$, as it already caches $V_{i}$ and $V_{i+1}$. Then, using the decoded symbols $V_{i-1}$ and $V_{i+2}$, node $n_{i}$ can then proceed to decode new symbols $V_{i-2}$ and $V_{i+3}$. This successive decoding process continues further for $V_{i-3}$ and $V_{i+4}$. In other words, each node decodes symbols sequentially, starting from those sent by nearby nodes and progressing to those from farther nodes. The decoding order is presented in Table II. The symbol ‘*’ at row $n_{i}$ and column $V_{j}$, where $i\in[8]$ and $j\in[8]$, indicates that node $n_{i}$ generates (and thus has) the IV $V_{j}$. In row $n_{i}$, integer $1$ indicates that the IV in the corresponding column can be directly decoded by node $n_{i}$ upon receiving the encoded symbol broadcast by its neighboring nodes. The integer $2$ signifies that decoding the corresponding IV depends on the symbol associated with the nearest integer $1$ in the same row. Similarly, the integer $3$ denotes that decoding the corresponding IV relies on the symbol associated with the nearest integer $2$ in the same row.
TABLE II: Decoding order for $\mathsf{N}=8$, $r=2$ and $d=3$
@@ IV-B General Scheme
If $d\geq r$, node $n_{i}$ can also decode $2d$ desired symbols but with successive decoding. Taking $Y_{i_{-}}^{(1)}$ as an example (illustrated in Fig. 9), the same color on the two sides of the summations means the different coded symbols share components. Since node $n_{i}$ has already known the right-hand side of the first r1r-1 summations (i.e., $V_{i+r-2},\ldots,V_{i}$), it can decode their left-hand side. These decoded symbols then serve as known components for the next r1r-1 summations (i.e., the $r^{\textnormal{th}}$ to $(2r-1)^{\textnormal{th}}$ summations from the bottom), allowing further decoding in a successive manner. Following this process, node $n_{i}$ successively decodes $2d$ desired symbols
@@
We now describe the transmission process from time clock $k\geq 2$. After time k1k-1, node $n_{i}$ has successfully decoded and stored the following symbols,
@@ V Coded Transmission for All-to-all Computing
For clarity, we redefine the notation for transmitted messages. The message $X_{i}^{(t)}$, transmitted by node $n_{i}$ at time slot $t$, is now written as $X_{i}^{(j,k)}$. Here, $X_{i}^{(j,k)}$ represents the message transmitted by node $n_{i}$ at the $k^{\textnormal{th}}$ step of round $j$, where $0\leq k\leq j$ and $t=1+2+\ldots+(j-1)+k$. Similarly, we re-write $Y_{i}^{(t)}$ as $Y_{i}^{(j,k)}$, maintaining the same relationship $0\leq k\leq j$ and $1+2+\ldots+(j-1)+k=t$.
@@ V-A An Illustrative Example
In the first round, each node transmits packets with distance $1$ to their destination nodes, i.e., $W_{i}^{(1)}$, $i\in[\mathsf{N}]$. The encoding and transmission process is presented in Table IV. For each node $i$, the coded packet $X_{i}^{(1,1)}=v_{i}^{i+1}\oplus v_{i+2}^{i-1}$ is broadcast to neighbor nodes $n_{i-1}$ and $n_{i+1}$. For example, node $n_{1}$ broadcasts $X_{1}^{(1,1)}=v_{1}^{2}\oplus v_{3}^{8}$, it also receives the encoded packets $X_{2}^{(1,1)}=v_{2}^{3}\oplus v_{4}^{1}$ and $X_{8}^{(1,1)}=v_{8}^{1}\oplus v_{2}^{7}$ which are broadcast by node $n_{2}$ and $n_{8}$, respectively. Since node $n_{1}$ has cached $v_{2}^{7}$ and $v_{2}^{3}$, it can decode the desired packets $v_{8}^{1}$ and $v_{4}^{1}$ by subtracting the known values from $X_{2}^{(1,1)}=v_{2}^{3}\oplus v_{4}^{1}$ and $X_{8}^{(1,1)}=v_{8}^{1}\oplus v_{2}^{7}$, respectively. The same decoding process applies to all other nodes.
In the second round, it will take $2$ steps of each node to transmit the packets that have a distance of $2$, i.e., $W_{i}^{(2)}$, $i\in[\mathsf{N}]$. The encoding and transmission process is shown in Table IV. For each node, the encoded packets of the first broadcast are generated from the packets in $W_{i}^{(2)}$. After receiving the coded packets from the broadcast in the first step, each node eliminates the cached content and re-encodes the remaining packets for the broadcast in the second step. For example, node $n_{1}$ received $X_{8}^{(2,1)}=v_{8}^{2}\oplus v_{2}^{6}$ from node $n_{8}$ and $X_{2}^{(2,1)}=v_{2}^{4}\oplus v_{4}^{8}$ from node $n_{2}$ in the first step. Node $n_{1}$ can decode $v_{8}^{2}$ and $v_{4}^{8}$ since it has cached $v_{2}^{6}$ and $v_{2}^{4}$. Then $n_{1}$ generates a new coded packet $X_{1}^{(2,2)}=v_{8}^{2}\oplus v_{4}^{8}$ and broadcasts it in the second step.
The third round requires $3$ steps to send the packets $W_{i}^{(3)}$, $i\in[\mathsf{N}]$, as shown in Tables V. The coding process follows the same principle as in previous rounds. However, we can observe that in this case, each packet in $W_{i}^{(3)}$ will be sent by two nodes. For example, $v_{1}^{4}$ appears in both coded packets $X_{1}^{(3,1)}=v_{1}^{4}\oplus v_{3}^{6}$ and $X_{7}^{(3,1)}=v_{7}^{2}\oplus v_{1}^{4}$, broadcast by $n_{1}$ and $n_{7}$, respectively. Therefore, the two nodes can each send half of the packets to save on transmission costs. Finally, the NCL can be calculated as $\mathsf{T}_{2}=1+2+2.5=5.5$. In this example, the decoding at each step relies solely on initially cached content. However, as we will see in the next section, from the third step onward, decoding also depends on the undesired symbols decoded in the previous two steps.
TABLE III: Round 11: Step 11.
Round 11 Step 11
@@
TABLE IV: Round 22: Step 11 and Step 22.
Round 22 Step 11 Step 22
@@
TABLE V: Round 33: Step 11, step $2$ and Step 33.
Round 33 Step 11 Step 22 Step 33
@@ V-B1 Proposed scheme for $d=1$
As described in Section V-A, the transmitted IVs of node $i$ can be categorized into different sets $W_{i}^{(1)},\ldots,W_{i}^{(\lceil\frac{\mathsf{N}-r}{2}\rceil)}$. Since the packets follow the cyclic placement, the desired packets are located at most $\left\lceil\frac{\mathsf{N}-r}{2}\right\rceil$ hops away from node $n_{i}$.
The transmissions are performed in rounds, with packets at different distances being sent in separate rounds.
@@
${=\frac{1}{2}\left(\left\lceil\frac{\mathsf{N}-r}{2}\right\rceil\right)^{2}+\frac{1}{2}\left\lceil\frac{\mathsf{N}-r}{2}\right\rceil}.$ (23)
@@ V-B2 Modification for $1<d\leq 2(r-1)$
The number of broadcasts can be further reduced when the nodes have a larger broadcast distance. Based on the transmission strategy for $d=1$, here is the modified version for $d\leq 2(r-1)$. Unlike the case of $d=1$ where the $3$ adjacent nodes perform reverse carpooling, here we allow nodes that are distance $d$ or r1r-1 apart to perform reverse carpooling. The paths of information flow are demonstrated in Fig. 11.
@@
$=\underbrace{1+\ldots+1}_{d_{1}}+\underbrace{2+\ldots+2}_{d}+\ldots+\!\underbrace{\left\lceil\frac{\left\lceil\frac{\mathsf{N}-r}{2}\right\rceil\!-\!d_{1}}{d}\right\rceil\!+\!1\!+\!\ldots\!+\!\left\lceil\frac{\left\lceil\frac{\mathsf{N}-r}{2}\right\rceil\!-\!d_{1}}{d}\right\rceil\!+\!1}_{\left(\lceil\frac{\mathsf{N}-r}{2}\rceil-d_{1}\right)_{\bmod d}}$
${=\frac{d}{2}\!\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil-d_{1}}{d}\right\rfloor^{2}\!+\!\frac{3d}{2}\!\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil\!-\!d_{1}}{d}\right\rfloor\!\!+\!d_{1}\!+\!\left(\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil\!-\!d_{1}}{d}\right\rfloor\!+\!2\right)\left(\left\lceil(\mathsf{N}-r)/2\right\rceil\!-\!d\right)_{d}}$
${=\left\{\begin{array}[]{cc}\vskip 5.0pt\frac{d}{2}\!\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil-r+1}{d}\right\rfloor^{2}\!+\!\frac{3d}{2}\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil-r+1}{d}\right\rfloor\!+\!r\!-\!1\!+\!\left(\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil-r+1}{d}\right\rfloor\!+\!2\right)\left(\left\lceil(\mathsf{N}-r)/2\right\rceil\!-\!d\right)_{d}&\begin{subarray}{c}d_{1}=r-1,\\ 1<r,\end{subarray}\\ \frac{d}{2}\!\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil}{d}\right\rfloor^{2}\!+\!\frac{d}{2}\!\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil}{d}\right\rfloor\!+\!\left(\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil}{d}\right\rfloor+1\right)\left(\left\lceil(\mathsf{N}-r)/2\right\rceil\!-\!d\right)_{d}&\begin{subarray}{c}d_{1}=d,\\ 1<r,\end{subarray}\end{array}\right.}$ (29)
where $(x)_{d}$ denotes $x\bmod d$ for $x\in\mathbb{Z}$.
@@ V-B3 Modification for $2r-1\leq d$
to nodes $n_{i-d}$ and $n_{i+d}$, respectively. For $i\in[\mathsf{N}],j\in[\lceil\frac{\mathsf{N}-r}{2}\rceil]$, $X_{i}^{(j,0)}$ constitute the additional step before the first regular transmission. These packets provide the necessary side information for decoding in the first step, enabling the pipeline to proceed according to the same decoding progression as described previously. Therefore, in this case, we can calculate the NCL as
@@
$\mathsf{T}_{\textnormal{2-cyc}}(r,d)$ $=\underbrace{1+\ldots+1}_{d}+\underbrace{2+\ldots+2}_{d}+\ldots+\!\underbrace{\left\lceil\frac{\left\lceil\frac{\mathsf{N}-r}{2}\right\rceil}{d}\right\rceil\!+\!\ldots\!+\!\left\lceil\frac{\left\lceil\frac{\mathsf{N}-r}{2}\right\rceil}{d}\right\rceil}_{\lceil\frac{\mathsf{N}-r}{2}\rceil_{\bmod d}}+\left\lceil\frac{\mathsf{N}-r}{2}\right\rceil$
@@
$={\frac{d}{2}\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil}{d}\right\rfloor^{2}+\frac{d}{2}\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil}{d}\right\rfloor+\left\lfloor\frac{\left\lceil(\mathsf{N}-r)/2\right\rceil}{d}\right\rfloor\left(\left\lceil(\mathsf{N}-r)/2\right\rceil\right)_{d}+\left\lceil(\mathsf{N}-r)/2\right\rceil}.$ (30)
@@
where $(x)_{d}$ denotes $x\bmod d$ for $x\in\mathbb{Z}$. From (23), (29) and (30), we can conclude that the achievable NCL of the proposed scheme is as stated in Theorem 2.
VI Concluding Remark and Future Directions